ADNAN MUHAMMED

Grade 12,

how can we calculate the relative velocity between rain aND man easily?

how can we calculate the relative velocity between rain aND man easily?

Grade:11

1 Answers

MIR ROUF
11 years ago
  1. In case 1        Vrp=Vr - Vp            
  2. In case 2       Vrp2 =Vr2-Vp2    Also Sin (angle)=Vp/Vr
  3. Vrp2=             Vr2+Vp2-2VrVpsin(angle),

                                            Tan (angle)=Vp-VrSin(angle)/VrCos(angle)